# MAROON: A Framework for the Joint Characterization of Near-Field High-Resolution Radar and Optical Depth Imaging Techniques

### v2

@@ 4.2.2. Time-Resolved Sensors (Time of Flight)
Transmitters Type -- Laser Projector LED Array TDM Antenna Array
Array Size -- -- -- 2$\times$47 TX $\updownarrow$
Wavelength -- 840–860 nm 860 nm 3.6-4.2 mm
Frequency -- $\approx$ 353 THz $\approx$ 353 THz 72-82 GHz
Signal Modulation -- Spatial Multiplexing AMCW FSCW ($\text{N}_{f}=128$)
@@
Array Size 2$\times$1928$\times$1208 px 2$\times$1280$\times$800 px 1024$\times$1024 px 2$\times$47 RX $\leftrightarrow$
Spatial Size* 2$\times$5.8$\times$3.6 mm 2$\times$3.8$\times$2.4 mm 3.6$\times$3.6 mm 2$\times$13.8 cm
Depth Image Resolution 1920$\times$1080 px 1280$\times$720 px 640$\times$576 px 301$\times$301 px
@@
Table 1. Overview of the sensors and their parameters used in our experiments. Rows with * indicate information not directly given by the manufacturer. $\delta_{x}$ and $\delta_{y}$ of camera-based systems is approximately determined from the per-pixel field of view. Spatial resolution formulae are provided in the supplementary material. Due to missing data for the Azure Kinect, $\delta_{z}$ is assumed to be theoretically higher than the depth accuracy given in (Bamji et al., 2018).
@@ 5. The MAROON Dataset
The capture of the MAROON dataset allows for a comprehensive analysis with respect to the characteristics of the four previously described depth sensing techniques. This section outlines the capture setup and data acquisition pipeline, depicted in Figure 1, to aid future research on the publicly accessible data. Four single-view depth sensors are used in our experiments: Stereolabs ZED X (Stereolabs, 2023) (Passive Stereo), Intel Realsense D435i (Intel, 2023) (Active Stereo), Microsoft Azure Kinect (Microsoft, 2022) (NIR ToF), and a submodule of Rohde & Schwarz’s QAR50 (Rohde & Schwarz, 2023) (RF ToF). Example data captured from these sensors is visualized in Figure 2.
@@ 5.1. Sensor Setup
Our sensor setup consists of four mounted single-view depth sensors and a ground truth (GT) optical multi-view stereo (MVS) system comprising five calibrated DSLR cameras, which are depicted on the left in Figure 1. While all single-view sensors are designed to achieve an optimal balance between depth quality and acquisition time, the MVS system employs an offline reconstruction process that is specifically optimized for depth quality. In summary, eight cameras are mounted on tripods and arranged around the MIMO imaging radar on a desk, thereby maximizing the area of intersection of each sensor’s field of view, to ensure similar object visibility. All sensors and the GT system are time-synchronized, either through hardware or software, to capture the object at the same moment.
Prior to capture, the object is positioned in the center of the squared radar aperture, and approximately at the center of the joint field-of-view intersection, propped up with boards crafted from styrofoam — a material that is considered to be nearly fully penetrated by the RF signal — to prevent external interference of radio waves from other sources in the vicinity, apart from the object of interest. For similar reasons, absorbers are placed behind the object of interest. Similarly, for optical sensors a loose black cloth, which is penetrated by RF signals, is suspended in front of the absorbers to visually occlude the room’s background. The sensor settings are chosen with respect to a trade-off between fair sensor comparability and practical applicability. More details can be found in the supplementary material. An overview of the chosen settings, together with relevant sensor parameters, is given in Table 1.
@@ 5.2. Data Acquisition Pipeline
To examine the characteristics of various objects in relation to different materials, geometries, and distances from the sensor, we recorded the MAROON dataset, which comprises static and quasi-static targets captured from all sensors simultaneously. With respect to the order of steps described in Figure 1 (middle), we will now continue to elaborate on the details of the acquisition pipeline.
@@
We spatially aligned the coordinate systems of each depth imager using the calibration method by Wirth et al. (2024). In this method, four respective spherical objects of styrofoam and metal, tailored to the visibility of optical and RF sensors, are captured. In the sensor-specific reconstructions, these spheres are automatically located and jointly aligned using spatial registration. This approach enables a direct comparison of the object reconstructions in a metrical space. Calibration errors are expected to be in 1–2 mm range with respect to the Chamfer distance, in analogy with Wirth et al. (2024).
The five DSLR cameras of the MVS system are treated as a unified sensor with a common coordinate system, which is spatially calibrated with that of the depth imagers. The camera extrinsic and intrinsic parameters of the MVS system are determined from images capturing a conventional optical calibration target with a checkerboard pattern. For this, we use the software provided by Agisoft Metashape. Remaining calibration errors exhibit a root mean square reprojection error of 0.38 px, averaged over all camera calibrations performed during the dataset capture.
@@
The reconstruction method of MVS is similar to passive stereo imagers. Hence, inaccurate reconstructions can be the result when dealing with textureless and view-dependent object materials. To circumvent this limitation, we implement a distinct capture process for particularly challenging objects to generate more reliable GT reconstructions. Once the object has been captured by all depth imagers, a thin multicolored speckle pattern is applied using water colors that assists the correspondence finding by a subsequent MVS-only capture. In order to ensure exact alignment between that GT reconstruction and other imaging modalities, the speckle is applied in situ without moving the object.
@@
To perform an accurate and precise object-centric sensor evaluation, it is essential to isolate the estimated object depth from the background. For optical systems, we acquire segmentation masks by performing a semi-automatic foreground-background segmentation. Given that all depth imagers capture RGB images — either for depth estimation or via a separate calibrated camera — we first segment the RGB images using manually defined object labels in conjunction with Grounded-SAM (Ren et al., 2024). This generates a binary segmentation mask of the object, $\bm{M}$, where all valid pixels $(u,v)$ are included in $\bm{M}^{+}(u,v)=\{\bm{M}(u,v)>0\}$. We then manually correct failure cases in the resulting segmentation masks. The same procedure is employed to MVS images to produce masked GT reconstructions. For the imaging radar, the voxel volume of the BP algorithm (Equation 7) is constrained to enclose only the object of interest. In this way, segmentation masks are automatically determined from the valid pixels remaining after depth estimation.
Depth Estimation
[…]
@@ 6. Evaluation
In this section, we compare the reconstructions produced by the four presented depth imagers with a ground-truth reconstruction in a common metric space and describe the methods used in this process. Subsequently, the results of these methods are presented.
6.1. Methods
In accordance with the order in Figure 1 (right), the necessary steps for evaluation are explained.
Reconstruction
We acquire a point cloud of the object’s surface utilizing the 2D depth and auxiliary data provided by MAROON. For a given pixel position $(u,v)$ and its corresponding depth $d$ from an optical depth sensor, we first verify its validity using the segmentation map of the same resolution — a step that has already been performed for radar during depth filtering. Subsequently we project each valid triple $(u,v,d)$ back into 3D space using the given transformation matrix $\bm{T}\in\mathbb{R}^{4\times 4}$:
[…]
For all optical depth imagers, this equation is the inverse of a perspective transformation with intrinsic camera matrix $\bm{I}\in\mathbb{R}^{3\times 3}$, pixel offset vector $\bm{t}\in\mathbb{R}^{3}=\bm{0}$ and $a=d$. Analogously for radar data, the equation is the inverse of an orthographic transformation with a scale matrix $\bm{I}$, pixel offset $\bm{t}$, and $a=1$.
During evaluation, we average valid depth values across 10 frames for static objects to incorporate temporal characteristics and reduce random noise. We do not average quasi-static objects (e.g., human hands) and instead take the first frame, as it is closest to the point in time where the GT captures have been taken.
Joint Alignment
To estimate the deviation of a sensor reconstruction $\bm{R}_{s}\in\mathbb{R}^{M\times 3}$ from the GT, $\bm{R}_{g}\in\mathbb{R}^{N\times 3}$, we use the previously determined spatial calibration parameters ${\bm{K}_{g\rightarrow s}\in\mathbb{R}^{4\times 4}}$ to transform $\bm{R}_{g}$ from the GT space $g$ into the sensor space $s$:
(10) $\widetilde{\bm{R}}_{g}^{s}=\widetilde{\bm{R}}_{g}\bm{K}_{g\rightarrow s}^{T}\;.$
$\widetilde{\bm{R}}$ denotes the homogeneous version of $\bm{R}$. We use the notation $\bm{R}^{*}$ to indicate a reconstruction that has been transformed to sensor space *.
Depth Deviation Measurement
For each object, we compute the point-wise deviation between a sensor and the GT reconstruction with respect to two metrics: one-sided Chamfer distance and one-sided projective error. The one-sided Chamfer distance, C, is computed per point $\bm{p}\in\mathbb{R}^{3}$ in the source point cloud $\bm{P}\in\mathbb{R}^{N\times 3}$ with respect to the distance to the nearest point $\bm{q}\in\mathbb{R}^{3}$ in the destination point cloud $\bm{Q}\in\mathbb{R}^{M\times 3}$:
@@ 6.1. Methods
For both, C and P, the respective subscripts $p$ and $u,v$ are used as placeholders to denote the points and pixels used in the metric computation. We compute both metric types for each reconstruction pair of a sensor, $\bm{R}_{s}$, and the GT, $\bm{R}_{g}$. Since the one-sided Chamfer distances are sensitive to the point cloud density, we uniformly sample the points from the sensor and the GT with respect to a common image pixel grid. We achieve this by, first, computing a simulated depth map $\widehat{\bm{D}}_{g}^{s}$ in the image space of the sensor $s$ and, second, reconstructing $\widehat{\bm{R}}^{s}_{g}$ from this depth map, using Equation 9 with the sensor-specific transformation $\bm{T}_{s}$. The simulated depth map is computed by rasterizing a triangulated representation of $\bm{R}^{s}_{g}$ with respect to $\bm{T}_{s}$. In this way, we also discard points in optical sensors $\bm{R}^{s}_{g}$ that are not visible in the view of camera $s$. The resulting depth map $\widehat{\bm{D}}_{g}^{s}$ is additionally used to measure the projective error. To summarize, we compute: $\text{C}_{\text{g}}$ Chamfer distance ground truth. $\forall\bm{g}\in\widehat{\bm{R}}^{s}_{g}:\text{C}_{\bm{g}}(\bm{R}_{s})$ $\text{C}_{\text{s}}$ Chamfer distance sensor. $\forall\bm{s}\in\bm{R}_{s}:\text{C}_{\bm{s}}(\widehat{\bm{R}}^{s}_{g})$ P Projective error. $\forall(u,v)\in\bm{M}^{+}(u,v):\text{P}_{u,v}(\bm{D}_{s},\widehat{\bm{D}}_{g}^{s})$ for ${\bm{M}=\bm{M}_{s}\cap\widehat{\bm{M}}_{g}^{s}}$, where $\bm{M}_{s}=\{\bm{D}_{s}>0\}$ and $\widehat{\bm{M}}_{g}^{s}=\{\widehat{\bm{D}}_{g}^{s}>0\}$ describe the intersection masks of valid pixels from the sensor and the projected GT, respectively. $\text{P}_{\text{e}}$ Projective error with erosion. $\forall(u,v)\in\bm{M}_{e}^{+}(u,v):\text{P}_{u,v}(\bm{D}_{s},\widehat{\bm{D}}_{g}^{s})$ for ${\bm{M}_{e}=\bm{M}_{s}\cap f(\widehat{\bm{M}}_{g}^{s})}$, where $f(\widehat{\bm{M}}_{g}^{s})$ is a function performing mask erosion using a kernel size of $K\times K$ pixels. The size $K\in[0,20]$ is semi-manually selected for each object and sensor, and included together with other evaluation metadata in the release of our dataset.
Metric Type Silhouette Noise Missing Surfaces 3D Error Depth Error
@@ 6.2. Results
In this section, we first present the evaluation results, quantified using four complementary metrics: $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, and $\text{P}_{\text{e}}$. Each metric is sensitive to different aspects, as detailed in Table 3. Important to note is that the computed quantities do not perfectly reflect the absolute depth error in the real world, due to the limitations of our experimental setup, which is constrained by the accuracy of the GT system and the spatial calibration. Instead, the results should be interpreted as depth deviation from the ground truth, referred to as sensor-GT deviation. We begin by presenting the sensor-GT deviation results in relation to various objects and different object-to-sensor distances. Given that near-field imaging radars are less explored compared to optical depth sensors, we dedicate the latter part of this section to the radio-frequency signal response. Finally, we note that in this section the results are objectively presented, reserving further interpretations for subsequent sections, where they will be discussed from a general perspective (Section 7), with specific attention to partially transmissive media (Section 8), and focusing on the RF signal response (Section 9).
6.2.1. Depth Deviation at Single and Multiple Object Distances

### v4

@@ 4.2.2. Time-Resolved Sensors (Time-of-Flight)
Depth Processing Time* ¡ 33 ms ¡ 33 ms ¡ 33 ms $\approx$ 78 s
Transmitters Type - Laser Projector LED Array TDM Antenna Array
[…]
Wavelength - 840–860 nm 860 nm 3.6-4.2 mm
Frequency - $\approx$ 353 THz $\approx$ 353 THz 72-82 GHz
Signal Modulation - Spatial Multiplexing AMCW FSCW ($\text{N}_{f}=128$)
@@
[…]
Field of view $110^{\circ}\times 80^{\circ}$ $87^{\circ}\times 58^{\circ}$ $75^{\circ}\times 65^{\circ}$ $\approx 53^{\circ}\times 53^{\circ}$ *
[…]
Spatial
Resolution*
$\bm{\delta_{x}\times\delta_{y}\times\delta_{z}}$
@@
Table 1. Overview of the sensors and their parameters used in our experiments. Rows with * indicate derived information not directly given by the manufacturer. Depth processing times were computed on a system with an NVIDIA GeForce RTX 3080 graphics card (10GB VRAM) and an Intel Xeon W-1390P (3.50 GHz) processor. Note that due to its fundamentally different operating principle, modeling the field of view of the QAR50 similar to a camera is a very simplified approximation, and we refer to the supplementary material for further details. $\delta_{x}$ and $\delta_{y}$ of camera-based systems is approximately determined from the per-pixel field of view. Spatial resolution formulae are provided in the supplementary material. Due to missing data for the Azure Kinect, $\delta_{z}$ is assumed to be theoretically higher than the depth accuracy given in (Bamji et al., 2018).
@@ 5. The MAROON Dataset
The capture of the MAROON dataset allows for a comprehensive analysis with respect to the characteristics of the four previously described depth sensing techniques.
To accomplish this, we collected a diverse set of common household and construction objects. We ensured having a broad variety of materials and geometries, with varying complexity, which we selected based on prior knowledge of the sensor operating principles (see Section 4.2). The selection aimed to identify challenging objects for reconstruction, highlighting the limitations of current depth imagers and providing a valuable data resource for improving these technologies, e.g. through integration of multimodal sensor data.
In the further course of this section, we outline the capture setup and data acquisition pipeline, depicted in Figure 1, to aid future research on the publicly accessible data. Four single-view depth sensors are used in our experiments: Stereolabs ZED X (Stereolabs, 2023) (Passive Stereo), Intel Realsense D435i (Intel, 2023) (Active Stereo), Microsoft Azure Kinect (Microsoft, 2022) (NIR ToF), and a submodule of Rohde & Schwarz’s QAR50 (Rohde & Schwarz, 2023) (RF ToF).
@@ 5.1. Sensor Setup
Our sensor setup consists of four mounted single-view depth sensors and a ground-truth (GT) optical multi-view stereo (MVS) system comprising five calibrated DSLR cameras, which are depicted on the left in Figure 1. While all single-view sensors are designed to achieve an optimal balance between depth quality and acquisition time, the MVS system employs an offline reconstruction process that is specifically optimized for depth quality. In summary, eight cameras are mounted on tripods and arranged around the MIMO imaging radar on a desk, thereby maximizing the area of intersection of each sensor’s field of view, to ensure similar object visibility. All sensors and the GT system are time-synchronized, either through hardware or software, to capture the object at the same moment.
Prior to capture, the object is positioned in the center of the squared radar aperture, and approximately at the center of the joint field-of-view intersection, propped up with boards crafted from styrofoam — a material that is considered to be nearly fully penetrated by the RF signal — to prevent external interference of radio waves from other sources in the vicinity, apart from the object of interest. For similar reasons, absorbers are placed behind the object of interest. Similarly, for optical sensors a loose black cloth, which is penetrated by RF signals, is suspended in front of the absorbers to visually occlude the room’s background. The sensor settings are chosen with respect to a trade-off between fair sensor comparability and practical applicability (see supp. mat.). An overview of the chosen settings, together with relevant sensor parameters, is given in Table 1.
@@ 5.2. Data Acquisition Pipeline
The MAROON dataset comprises static and quasi-static, i.e., with slow, minimal motion as in case for human hands, targets of differing materials and geometries, captured at multiple distances from all sensors simultaneously. With respect to the order of steps described in Figure 1 (middle), we will now continue to elaborate on the details of the acquisition pipeline.
@@
We spatially aligned the coordinate systems of each depth imager using the calibration method by Wirth et al. (2024). In this method, four respective spherical objects of styrofoam and metal, tailored to the visibility of optical and RF sensors, are captured. In the sensor-specific reconstructions, these spheres are automatically located and jointly aligned using spatial registration. This approach enables a direct comparison of the object reconstructions in a metrical space. Calibration errors are expected to be in 1–2 mm range with respect to the Chamfer distance, in analogy with the evaluation scheme used in Wirth et al. (2024).
The five DSLR cameras of the MVS system are treated as a unified sensor with a common coordinate system, which is spatially calibrated with that of the depth imagers. The camera extrinsic and intrinsic parameters of the MVS system are determined from images capturing a conventional optical calibration target with a checkerboard pattern. For this, we use the commercial software provided by Agisoft Metashape. Remaining calibration errors exhibit a root mean square reprojection error of 0.38 px, averaged over all camera calibrations performed during the dataset capture.
@@
The reconstruction method of MVS is similar to passive stereo imagers. Hence, inaccurate reconstructions can be the result when dealing with textureless and view-dependent object materials. To circumvent this limitation, we implement a distinct capture process for a subset of particularly challenging objects to generate more reliable GT reconstructions. After the object has been captured once by all sensors (including MVS), a thin multicolored speckle pattern is applied using water colors that assists the correspondence finding of the subsequent, additional MVS-only capture. In order to ensure exact alignment between that GT reconstruction and other imaging modalities, the speckle is applied in situ without moving the object.
@@
To perform an accurate object-centric sensor evaluation, it is essential to isolate the estimated object depth from the background. For optical systems, we acquire segmentation masks by performing a semi-automatic foreground-background segmentation. Given that all depth imagers capture RGB images — either for depth estimation or via a separate calibrated camera — we first segment the RGB images using manually defined object labels in conjunction with Grounded-SAM (Ren et al., 2024). This generates a binary segmentation mask of the object, $\bm{M}$, where all valid pixels $(u,v)$ are included in $\bm{M}^{+}(u,v)=\{\bm{M}(u,v)>0\}$. We then manually correct failure cases in the resulting segmentation masks. The same procedure is employed to MVS images to produce masked GT reconstructions. For the imaging radar, the voxel volume of the BP algorithm (Equation 7) is constrained to enclose only the object of interest. In this way, segmentation masks are automatically determined from the valid pixels remaining after depth estimation.
Reconstruction and Depth Estimation
MAROON offers raw sensor data, along with intermediate and final reconstruction output, stored in various data representations depending on each depth imager.
For optical depth sensors, we store RGB images, auxiliary data such as infrared measurements, and depth maps, which are obtained using the corresponding signal processing algorithms provided by the manufacturer.
The imaging radar captures raw measurements in form of a tensor of $N_{\text{RX}}\times N_{\text{TX}}\times N_{f}$ complex numbers, where $N_{\text{RX}}=N_{\text{TX}}=94$ and $N_{f}=128$. They are stored alongside the volumetric output produced after backprojection, as well as post-processed 2D depth and confidence maps. Using the raw tensor, we perform the BP algorithm on a $301\times 301\times 201$ voxel grid, with voxel centers uniformly sampled within a $30\times 30\times 20$ $\text{cm}^{3}$ volume around the object center, yielding volumetric data that is stored as intermediate output. Subsequently, we apply maximum projection (Equation 8) to acquire a 2D projection of the depth as well as a 2D confidence map. Using the latter, we filter out depth values according to a threshold of 14-14 dB relative to the maximum value. As mentioned in Section 4.2.2, such thresholding is challenging for unknown objects. We chose this threshold empirically over all objects in the dataset, aiming at a good balance between pruning of noise and retention of object details, and provide an ablation study with different thresholds in the supplementary material. We encourage interested readers to experiment with different thresholds, using the raw radar data available in our dataset. After thresholding, the filtered result is stored as an orthographic depth map.
The ground-truth MVS setup captures five RGB images, which are stored alongside post-processed depth images and a mesh representation of the object, after performing reconstruction using Agisoft Metashape. Metashape (formerly Photoscan) commonly has a reconstruction accuracy in sub-millimeter range for similar capture environments (Mousavi et al., 2018; Remondino et al., 2014). After reconstruction, we finally apply Laplacian smoothing.
@@ 6. Evaluation
In this section, we compare the reconstructions produced by the four presented depth imagers with a ground-truth reconstruction in a common metric space and describe the metrics used in this process. Subsequently, the results of these methods are presented.
We note that in this section the results are objectively presented, reserving further interpretations for Section 7, where they will be discussed with specific attention to partially transmissive media (Section 7.2) and focusing on the RF signal response (Section 7.3).
6.1. Metrics
First, we average valid depth values of each sensor across 10 frames for static objects to incorporate temporal characteristics and reduce random noise. We do not average quasi-static objects, of which their reconstruction did not require the application of speckles, and instead take the first frame, as it is closest to the point in time where the GT captures without the speckle pattern have been taken.
Using the extrinsic calibration parameters (see supp. mat.), we subsequently transform the masked GT reconstruction, $\bm{R}_{g}$, into each sensor space $s$, yielding $\bm{R}_{g}^{s}$. We use the notation $\bm{R}^{*}$ to indicate a transformation to sensor space *.
Next, for each object, we compute the point-wise deviation between a sensor and the transformed GT reconstruction with respect to two metrics: one-sided Chamfer distance and one-sided projective error. The one-sided Chamfer distance, C, is computed per point $\bm{p}\in\mathbb{R}^{3}$ in the source point cloud $\bm{P}\in\mathbb{R}^{N\times 3}$ with respect to the distance to the nearest point $\bm{q}\in\mathbb{R}^{3}$ in the destination point cloud $\bm{Q}\in\mathbb{R}^{M\times 3}$:
@@ 6.1. Metrics
For both, C and P, the respective subscripts $\bm{p}$ and $u,v$ are used as placeholders to denote the points and pixels used in the metric computation.
Since the one-sided Chamfer distances are sensitive to the point cloud density, we uniformly sample the points from the sensor and the GT with respect to a common image pixel grid. We achieve this by, first, computing a simulated depth map $\widehat{\bm{D}}_{g}^{s}$ in the image space of the sensor $s$ and, second, reconstructing $\widehat{\bm{R}}^{s}_{g}$ from this depth map, using the inverse camera parameters to project it into 3D (see supp. mat.). The simulated depth map is computed by rasterizing a triangulated representation of $\bm{R}^{s}_{g}$ with respect to $\bm{T}_{s}$. In this way, we also discard points in optical sensors $\bm{R}^{s}_{g}$ that are not visible in the view of $s$. The resulting depth map $\widehat{\bm{D}}_{g}^{s}$ is additionally used to measure the projective error. To summarize, we compute: $\text{C}_{\text{g}}$ Chamfer distance ground truth. $\forall\bm{g}\in\widehat{\bm{R}}^{s}_{g}:\text{C}_{\bm{g}}(\bm{R}_{s})$ $\text{C}_{\text{s}}$ Chamfer distance sensor. $\forall\bm{s}\in\bm{R}_{s}:\text{C}_{\bm{s}}(\widehat{\bm{R}}^{s}_{g})$ P Projective error. $\forall(u,v)\in\bm{M}^{+}(u,v):\text{P}_{u,v}(\bm{D}_{s},\widehat{\bm{D}}_{g}^{s})$ for $\bm{M}=\bm{M}_{s}\cap\widehat{\bm{M}}_{g}^{s}$}, where $\bm{M}_{s}=\{\bm{D}_{s}>0\}$ and $\widehat{\bm{M}}_{g}^{s}=\{\widehat{\bm{D}}_{g}^{s}>0\}$ describe the intersection masks of valid pixels from the sensor and the projected GT, respectively. $\text{P}_{\text{e}}$ Projective error with erosion. $\forall(u,v)\in\bm{M}_{e}^{+}(u,v):\text{P}_{u,v}(\bm{D}_{s},\widehat{\bm{D}}_{g}^{s})$ for $\bm{M}_{e}=\bm{M}_{s}\cap f(\widehat{\bm{M}}_{g}^{s})$}, where $f(\widehat{\bm{M}}_{g}^{s})$ is a function performing mask erosion using a kernel size of $K\times K$ pixels. The size $K\in[0,20]$ is semi-manually selected for each object and sensor, and included together with other evaluation metadata in the release of our dataset.
Metric
Type
Silhouette
Noise
Missing
Surfaces
3D Error
Depth Error
@@ 6.2. Results
In this section, we first present the evaluation results, quantified using four complementary metrics: $\text{C}_{\text{g}}$, $\text{C}_{\text{s}}$, P, and $\text{P}_{\text{e}}$. Each metric is sensitive to different aspects, as detailed in Table 3. We begin by presenting the depth deviation in relation to various objects and different object-to-sensor distances. Given that near-field imaging radars are less explored compared to optical depth sensors, we dedicate the latter part of this section to the radio-frequency signal response.
6.2.1. Depth Deviation